Please help with this problem
Solve the linear systems by any method.
Eq 1 states: 1/3x - 3/4y = -1/2
Eq 2 states: 1/6x + 1/8y = 3/4

Answer by stanbon(75887)

Solve the linear systems by any method.
Eq 1 states: (1/3)x - (3/4)y = -1/2
Eq 2 states: (1/6)x + (1/8)y = 3/4
-------------------------------------
Multiply thru 1st by 12; Multiply thru 2nd by 24:
1st: 4x - 9y = -6
2nd: 4x + 3y = 18
----------------------
Subtract 2nd from 1st to get:
-12y = -24
y = 2
-----------------
Substitute into 1st to solve for x:
4x - 9*2 = -6
4x = 12
x = 3
--------------
Check in 2nd:
4*3 + 3*2 = 18
18= 18
=========================
Cheers,
Stan H.